OpenStudy (grace111):

Will medal Is 11/20 less or greater than 1/2? Is 0.21 less or greater than 1/2?

OpenStudy (anonymous):

how can we make 11/20 equal to 1/2?

OpenStudy (anonymous):

sorry let me rephrase that: what is the closest fraction to 11/20 that is equal to 1/2?

OpenStudy (anonymous):

perhaps 10/20? that being said, we had to subtract 1/20 from the equation to get it. Subtraction means its bigger than 1/2

OpenStudy (anonymous):

for the decimal, lets change it to a fraction. \[\frac{ 21 }{ 100 }\] now we change 1/2 to be equivalant, which means we just put 100 in the denominator. \[\frac{ 1 }{2}\] \[\frac{ 1 * 50 }{2 * 50}\] \[\frac{ 50 }{100}\] now we can obviously see which is greater, 21 or 50? :)

OpenStudy (grace111):

50 ?

OpenStudy (anonymous):

yep! so 1/2 is larger than the decimal :)